If y represents a number, which equation is a correct translation of the sentence?
20 subtracted from 7 times a number is 6.

Found 4 solutions by math_helper, Edwin McCravy, ikleyn, greenestamps:
Answer by math_helper(2461)
Answer by Edwin McCravy(20055)

He just blurted the answer out.  Here is what you do:

20 subtracted from 7 times a number is 6.
Replace the words "a number" by N:

20 subtracted from 7 times N is 6.
Replace the words "7 times N" by "7N":

20 subtracted from 7N is 6.
Replace the word "is" by an equal sign " = ":

20 subtracted from 7N = 6.
Remember that to subtract the first number from the second number,
we start with the second number and then we subtract the first number.
The first number is 20 and the second number is 7N.  So we start with the
second number 7N and then we subtract the first number 20 from it:

7N - 20 = 6 <--answer
Edwin

Answer by ikleyn(52781)

.

In my experience, explanations practically do not help in this kind of problems.

In this kind of problems, examples work much better, than explanations.



Answer by greenestamps(13200)



This is actually a bad example, because it is open to (at least) two different interpretations.

In the statement there are three things:
(1) numbers or expressions
(2) mathematical operations
(3) an equals sign

The task is to break the sentence down into those three kinds of things.

If read or spoken normally, this sentence breaks down into

(20) (subtracted from) (7 times a number) (is) (6)

So if y is the number, 7 times the number is 7y; and 20 SUBTRACTED FROM 7y means you start with the 7y and subtract 20: 7y-20.

The "is" is clearly the equals sign, so the mathematical equation is 7y-20 = 6.

However, this sentence is open to another very different, although far less likely, interpretation.

In written form, you could add commas to change the meaning:

"20, subtracted from 7, times a number, is 6"

In spoken form, you could accomplish the same thing using pauses in the reading:

"20 [pause] subtracted from 7 [pause] times a number [pause] is 6".

In that interpretation, 20 is subtracted from 7; the result is multiplied by the number, and the final result is 6.

Breaking the sentence into the three things as before, this gives us

(20) (subtracted from) (7) (times) (a number) (is) (6):

(7-20)*y = 6